OpenStudy (anonymous):

Not a big fan of trigonometric equations. Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. -tan^2x+sec^2x=sec^2x cos^2x

OpenStudy (anonymous):

sec^2x = 1/cos^2x

OpenStudy (anonymous):

and tan^2x = sin^2x over cos^2x

OpenStudy (anonymous):

sub them in and juggle around with the sin cos .. u shud can do it

OpenStudy (anonymous):

or u could just divide the both Left hand side and right hand side by sec^2x then u will get -tan^2x/sec^2x = cos^2x

OpenStudy (anonymous):

then sub in 1/cos^2x for sec^2x

OpenStudy (anonymous):